\begin{thebibliography}{99}

%\cite{TheMEG:2016wtm}
\bibitem{TheMEG:2016wtm}
  A.~M.~Baldini {\it et al.} [MEG Collaboration],\\
  %``Search for the lepton flavour violating decay $\mu ^+ \rightarrow \mathrm {e}^+ \gamma $ with the full dataset of the MEG experiment,''
  Eur.\ Phys.\ J.\ C {\bf 76} (2016) no.8,  434
%  doi:10.1140/epjc/s10052-016-4271-x
%  [arXiv:1605.05081 [hep-ex]].
  %%CITATION = doi:10.1140/epjc/s10052-016-4271-x;%%
  %103 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Baldini:2013ke}
\bibitem{Baldini:2013ke}
  A.~M.~Baldini {\it et al.},\\
  %``MEG Upgrade Proposal,''
  arXiv:1301.7225 [physics.ins-det].
  %%CITATION = ARXIV:1301.7225;%%
  %205 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Grange:2015fou}
\bibitem{Grange:2015fou}
  J.~Grange {\it et al.} [Muon g-2 Collaboration],\\
  %``Muon (g-2) Technical Design Report,''
  arXiv:1501.06858 [physics.ins-det].
  %%CITATION = ARXIV:1501.06858;%%
  %87 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Peccei:2006as}
\bibitem{Peccei:2006as}
  R.~D.~Peccei,\\
  %``The Strong CP problem and axions,''
  Lect.\ Notes Phys.\  {\bf 741} (2008) 3
%  doi:10.1007/978-3-540-73518-2_1
%  [hep-ph/0607268].
  %%CITATION = doi:10.1007/978-3-540-73518-2_1;%%
  %203 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Dobrescu:2014fca}
\bibitem{Dobrescu:2014fca}
  B.~A.~Dobrescu and C.~Frugiuele,\\
  %``Hidden GeV-scale interactions of quarks,''
  Phys.\ Rev.\ Lett.\  {\bf 113} (2014) 061801
%  doi:10.1103/PhysRevLett.113.061801
%  [arXiv:1404.3947 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevLett.113.061801;%%
  %32 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Dobrescu:2014ita}
\bibitem{Dobrescu:2014ita}
  B.~A.~Dobrescu and C.~Frugiuele,\\
  %``GeV-Scale Dark Matter: Production at the Main Injector,''
  JHEP {\bf 1502} (2015) 019
%  doi:10.1007/JHEP02(2015)019
%  [arXiv:1410.1566 [hep-ph]].
  %%CITATION = doi:10.1007/JHEP02(2015)019;%%
  %13 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Flacke:2016szy}
\bibitem{Flacke:2016szy}
  T.~Flacke, C.~Frugiuele, E.~Fuchs, R.~S.~Gupta and G.~Perez,\\
  %``Phenomenology of relaxion-Higgs mixing,''
  JHEP {\bf 1706} (2017) 050
%  doi:10.1007/JHEP06(2017)050
%  [arXiv:1610.02025 [hep-ph]].
  %%CITATION = doi:10.1007/JHEP06(2017)050;%%
  %15 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Lees:2017lec}
\bibitem{Lees:2017lec}
  J.~P.~Lees {\it et al.} [BaBar Collaboration],\\
  %``Search for invisible decays of a dark photon produced in e+e- collisions at BaBar,''
  arXiv:1702.03327 [hep-ex].
  %%CITATION = ARXIV:1702.03327;%%
  %9 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Lazzeroni:2017fza}
\bibitem{Lazzeroni:2017fza}
  C.~Lazzeroni {\it et al.} [NA62 Collaboration],\\
  %``Search for Heavy Neutrinos in $K^+ \rightarrow \mu^+ \nu_{\mu}$ Decays,''
  arXiv:1705.07510 [hep-ex].
  %%CITATION = ARXIV:1705.07510;%%
  %1 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Raggi:2014zpa}
\bibitem{Raggi:2014zpa}
  M.~Raggi and V.~Kozhuharov,\\
  %``Proposal to Search for a Dark Photon in Positron on Target Collisions at DA$\Phi$NE Linac,''
  Adv.\ High Energy Phys.\  {\bf 2014} (2014) 959802
%  doi:10.1155/2014/959802
%  [arXiv:1403.3041 [physics.ins-det]].
  %%CITATION = doi:10.1155/2014/959802;%%
  %25 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Bellgardt:1987du}
\bibitem{Bellgardt:1987du}
  U.~Bellgardt {\it et al.} [SINDRUM Collaboration],\\
  %``Search for the Decay mu+ ---> e+ e+ e-,''
  Nucl.\ Phys.\ B {\bf 299} (1988) 1.
%  doi:10.1016/0550-3213(88)90462-2
  %%CITATION = doi:10.1016/0550-3213(88)90462-2;%%
  %533 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Bertl:2006up}
\bibitem{Bertl:2006up}
  W.~H.~Bertl {\it et al.} [SINDRUM II Collaboration],\\
  %``A Search for muon to electron conversion in muonic gold,''
  Eur.\ Phys.\ J.\ C {\bf 47} (2006) 337.
%  doi:10.1140/epjc/s2006-02582-x
  %%CITATION = doi:10.1140/epjc/s2006-02582-x;%%
  %313 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Fael:2015gua}
\bibitem{Fael:2015gua}
  M.~Fael, L.~Mercolli and M.~Passera,\\
  %``Radiative $\mu$ and $\tau$ leptonic decays at NLO,''
  JHEP {\bf 1507} (2015) 153
%  doi:10.1007/JHEP07(2015)153
%  [arXiv:1506.03416 [hep-ph]].
  %%CITATION = doi:10.1007/JHEP07(2015)153;%%
  %16 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Fael:2016hnz}
\bibitem{Fael:2016hnz}
  M.~Fael and M.~Passera,\\
  %``Precision tests via radiative $\mu$ and $\tau$ leptonic decays,''
  arXiv:1602.00457 [hep-ph].
  %%CITATION = ARXIV:1602.00457;%%
  %4 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Pruna:2017upz}
\bibitem{Pruna:2017upz}
  G.~M.~Pruna, A.~Signer and Y.~Ulrich,\\
  %``Fully differential NLO predictions for the radiative decay of muons and taus,''
  arXiv:1705.03782 [hep-ph].
  %%CITATION = ARXIV:1705.03782;%%
  %1 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Flores-Tlalpa:2015vga}
\bibitem{Flores-Tlalpa:2015vga}
  A.~Flores-Tlalpa, G.~López Castro and P.~Roig,\\
  %``Five-body leptonic decays of muon and tau leptons,''
  JHEP {\bf 1604} (2016) 185
%  doi:10.1007/JHEP04(2016)185
%  [arXiv:1508.01822 [hep-ph]].
  %%CITATION = doi:10.1007/JHEP04(2016)185;%%
  %10 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Fael:2016yle}
\bibitem{Fael:2016yle}
  M.~Fael and C.~Greub,\\
  %``Next-to-leading order prediction for the decay $ \mu \to e\kern0.22em \left({e}^{+}{e}^{-}\right)\;\nu \kern0.2em \overline{\nu} $,''
  JHEP {\bf 1701} (2017) 084
%  doi:10.1007/JHEP01(2017)084
%  [arXiv:1611.03726 [hep-ph]].
  %%CITATION = doi:10.1007/JHEP01(2017)084;%%
  %3 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Pruna:2016spf}
\bibitem{Pruna:2016spf}
  G.~M.~Pruna, A.~Signer and Y.~Ulrich,\\
  %``Fully differential NLO predictions for the rare muon decay,''
  Phys.\ Lett.\ B {\bf 765} (2017) 280
%  doi:10.1016/j.physletb.2016.12.039
%  [arXiv:1611.03617 [hep-ph]].
  %%CITATION = doi:10.1016/j.physletb.2016.12.039;%%
  %3 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Blondel:2013ia}
\bibitem{Blondel:2013ia}
  A.~Blondel {\it et al.},\\
  %``Research Proposal for an Experiment to Search for the Decay $\mu \to eee$,''
  arXiv:1301.6113 [physics.ins-det].
  %%CITATION = ARXIV:1301.6113;%%
  %133 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Pruna:2014asa}
\bibitem{Pruna:2014asa}
  G.~M.~Pruna and A.~Signer,\\
  %``The $\mu\to e\gamma$ decay in a systematic effective field theory approach with dimension 6 operators,''
  JHEP {\bf 1410} (2014) 014
%  doi:10.1007/JHEP10(2014)014
%  [arXiv:1408.3565 [hep-ph]].
  %%CITATION = doi:10.1007/JHEP10(2014)014;%%
  %27 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Pruna:2015jhf}
\bibitem{Pruna:2015jhf}
  G.~M.~Pruna and A.~Signer,\\
  %``Lepton-flavour violating decays in theories with dimension 6 operators,''
  EPJ Web Conf.\  {\bf 118} (2016) 01031
%  doi:10.1051/epjconf/201611801031
%  [arXiv:1511.04421 [hep-ph]].
  %%CITATION = doi:10.1051/epjconf/201611801031;%%
  %7 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Crivellin:2016ebg}
\bibitem{Crivellin:2016ebg}
  A.~Crivellin, S.~Davidson, G.~M.~Pruna and A.~Signer,
  %``Complementarity in lepton-flavour violating muon decay experiments,''
  arXiv:1611.03409 [hep-ph].
  %%CITATION = ARXIV:1611.03409;%%
  %3 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Crivellin:2017rmk}
\bibitem{Crivellin:2017rmk}
  A.~Crivellin, S.~Davidson, G.~M.~Pruna and A.~Signer,\\
  %``Renormalisation-group improved analysis of $\mu\to e$ processes in a systematic effective-field-theory approach,''
  JHEP {\bf 1705} (2017) 117
%  doi:10.1007/JHEP05(2017)117
%  [arXiv:1702.03020 [hep-ph]].
  %%CITATION = doi:10.1007/JHEP05(2017)117;%%
  %2 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Bennett:2006fi}
\bibitem{Bennett:2006fi}
  G.~W.~Bennett {\it et al.} [Muon g-2 Collaboration],\\
  %``Final Report of the Muon E821 Anomalous Magnetic Moment Measurement at BNL,''
  Phys.\ Rev.\ D {\bf 73} (2006) 072003
%  doi:10.1103/PhysRevD.73.072003
%  [hep-ex/0602035].
  %%CITATION = doi:10.1103/PhysRevD.73.072003;%%
  %1408 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Hagiwara:2011af}
\bibitem{Hagiwara:2011af}
  K.~Hagiwara, R.~Liao, A.~D.~Martin, D.~Nomura and T.~Teubner,\\
  %``(g-2)_mu and alpha(M_Z^2) re-evaluated using new precise data,''
  J.\ Phys.\ G {\bf 38} (2011) 085003
%  doi:10.1088/0954-3899/38/8/085003
%  [arXiv:1105.3149 [hep-ph]].
  %%CITATION = doi:10.1088/0954-3899/38/8/085003;%%
  %525 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Hagiwara:2017zod}
\bibitem{Hagiwara:2017zod}
  K.~Hagiwara, A.~Keshavarzi, A.~D.~Martin, D.~Nomura and T.~Teubner,\\
  %``g − 2 of the muon: status report,''
  Nucl.\ Part.\ Phys.\ Proc.\  {\bf 287-288} (2017) 33.
%  doi:10.1016/j.nuclphysbps.2017.03.039
  %%CITATION = doi:10.1016/j.nuclphysbps.2017.03.039;%%

%\cite{Aoyama:2012wk}
\bibitem{Aoyama:2012wk}
  T.~Aoyama, M.~Hayakawa, T.~Kinoshita and M.~Nio,\\
  %``Complete Tenth-Order QED Contribution to the Muon g-2,''
  Phys.\ Rev.\ Lett.\  {\bf 109} (2012) 111808
%  doi:10.1103/PhysRevLett.109.111808
%  [arXiv:1205.5370 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevLett.109.111808;%%
  %201 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Gnendiger:2013pva}
\bibitem{Gnendiger:2013pva}
  C.~Gnendiger, D.~Stöckinger and H.~Stöckinger-Kim,\\
  %``The electroweak contributions to $(g-2)_\mu$ after the Higgs boson mass measurement,''
  Phys.\ Rev.\ D {\bf 88} (2013) 053005
%  doi:10.1103/PhysRevD.88.053005
%  [arXiv:1306.5546 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.88.053005;%%
  %97 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Jegerlehner:2009ry}
\bibitem{Jegerlehner:2009ry}
  F.~Jegerlehner and A.~Nyffeler,\\
  %``The Muon g-2,''
  Phys.\ Rept.\  {\bf 477} (2009) 1
%  doi:10.1016/j.physrep.2009.04.003
%  [arXiv:0902.3360 [hep-ph]].
  %%CITATION = doi:10.1016/j.physrep.2009.04.003;%%
  %666 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Colangelo:2014qya}
\bibitem{Colangelo:2014qya}
  G.~Colangelo, M.~Hoferichter, A.~Nyffeler, M.~Passera and P.~Stoffer,\\
  %``Remarks on higher-order hadronic corrections to the muon g−2,''
  Phys.\ Lett.\ B {\bf 735} (2014) 90
%  doi:10.1016/j.physletb.2014.06.012
%  [arXiv:1403.7512 [hep-ph]].
  %%CITATION = doi:10.1016/j.physletb.2014.06.012;%%
  %60 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Davier:2010nc}
\bibitem{Davier:2010nc}
  M.~Davier, A.~Hoecker, B.~Malaescu and Z.~Zhang,\\
  %``Reevaluation of the Hadronic Contributions to the Muon g-2 and to alpha(MZ),''
  Eur.\ Phys.\ J.\ C {\bf 71} (2011) 1515
   Erratum: [Eur.\ Phys.\ J.\ C {\bf 72} (2012) 1874]
%  doi:10.1140/epjc/s10052-012-1874-8, 10.1140/epjc/s10052-010-1515-z
%  [arXiv:1010.4180 [hep-ph]].
  %%CITATION = doi:10.1140/epjc/s10052-012-1874-8, 10.1140/epjc/s10052-010-1515-z;%%
  %632 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Kurz:2014wya}
\bibitem{Kurz:2014wya}
  A.~Kurz, T.~Liu, P.~Marquard and M.~Steinhauser,\\
  %``Hadronic contribution to the muon anomalous magnetic moment to next-to-next-to-leading order,''
  Phys.\ Lett.\ B {\bf 734} (2014) 144
%  doi:10.1016/j.physletb.2014.05.043
%  [arXiv:1403.6400 [hep-ph]].
  %%CITATION = doi:10.1016/j.physletb.2014.05.043;%%
  %73 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Calame:2015fva}
\bibitem{Calame:2015fva}
  C.~M.~Carloni Calame, M.~Passera, L.~Trentadue and G.~Venanzoni,\\
  %``A new approach to evaluate the leading hadronic corrections to the muon $g$-2,''
  Phys.\ Lett.\ B {\bf 746} (2015) 325
%  doi:10.1016/j.physletb.2015.05.020
%  [arXiv:1504.02228 [hep-ph]].
  %%CITATION = doi:10.1016/j.physletb.2015.05.020;%%
  %17 citations counted in INSPIRE as of 28 Jun 2017

%\cite{Abbiendi:2016xup}
\bibitem{Abbiendi:2016xup}
  G.~Abbiendi {\it et al.},\\
  %``Measuring the leading hadronic contribution to the muon g-2 via $\mu e$ scattering,''
  Eur.\ Phys.\ J.\ C {\bf 77} (2017) no.3,  139
%  doi:10.1140/epjc/s10052-017-4633-z
%  [arXiv:1609.08987 [hep-ex]].
  %%CITATION = doi:10.1140/epjc/s10052-017-4633-z;%%
  %7 citations counted in INSPIRE as of 28 Jun 2017

\end{thebibliography}

\documentclass[12pt,a4paper]{article}
\usepackage{amsmath,amssymb,epsfig,a4,latexsym}
\usepackage{physics}
\usepackage{slashed}
\usepackage{color}
\usepackage{rotating}
\usepackage{graphicx}
\usepackage{caption}
\usepackage[caption=false]{subfig}

\begin{document}
\thispagestyle{empty}
\begin{flushright}
PSI-PR-17-12\\
\today\\
\end{flushright}
\vspace{3em}
\begin{center}
{\large\bf The Standard Model and low-energy experiments: \\
from lepton-flavour violation to dark photons} 
\\ 
\vspace{3em}
{\sc G.~M.~Pruna\footnote{e-mail: Giovanni-Marco.Pruna@psi.ch}
}\\[2em]
{\sl Paul Scherrer Institut,\\
CH-5232 Villigen PSI, Switzerland}
\setcounter{footnote}{0}
\end{center}
\vspace{2ex}
\begin{abstract}
\noindent
This note is a concise theoretical summary of the first session on new physics searches at the high-intensity frontier of the IFAE2017 Conference.
Recent theoretical developments related to muonic lepton-flavour violation and $g-2$ are reviewed.
\end{abstract}
 
\newpage

\section*{Introduction}
\noindent
The Standard Model (SM) of particle physics with its most accurate description of fundamental interactions represents one of the greatest intellectual achievements of humankind. However, it does not satisfactorily explain the origin of matter, the nature of neutrino oscillations, the observation of dark matter and dark energy, and it does not accommodate gravity. Consequently, the SM is commonly accepted simply as a low-energy manifestation of an ultimate theory defined at the Planck energy scale, which should incorporate solutions to these open problems. The lack of evidence of any new physics (NP) signal from low- and high-energy experiments suggests that such NP completion could either be \emph{weakly coupled} to the SM or \emph{much heavier} than the electroweak (EW) symmetry-breaking scale (or a combination of both circumstances).

The experimental high-intensity frontier represents a very promising place to look for such NP scenarios: on the one hand, if beyond the SM (BSM) physics is realised at higher scales, then it is possible to search for deviations from the expected SM interactions (typically adopting an effective field theory (EFT) approach); on the other hand, lighter dark sectors can be investigated in searches for invisible decays and displaced interaction points.

During the first session on NP searches at the high-intensity frontier of the IFAE2017 Conference, the current~\cite{TheMEG:2016wtm} and future~\cite{Baldini:2013ke} experimental searches on muonic lepton-flavour violation, the status of the Muon $(g-2)$ experiment~\cite{Grange:2015fou}, the searches for invisible and/or exotic particles~\cite{Peccei:2006as,Dobrescu:2014fca,Dobrescu:2014ita,Flacke:2016szy} at BaBar~\cite{Lees:2017lec}, NA62~\cite{Lazzeroni:2017fza} and PADME~\cite{Raggi:2014zpa} were discussed. In what follows, we will focus on muonic experiments and briefly review some literature on the topic.

\section*{Testing the muonic sector of the SM}
\subsection*{Muonic lepton-flavour violation}
\noindent
To test efficiently for the presence of NP interactions, the experimental community has devoted a large effort to investigating processes that are forbidden/suppressed in the SM, due to the advantageous absence/suppression of background. One such promising study is the test of muonic lepton-flavour violating (LFV) decays, \emph{e.g.} $\mu\to e\gamma$~\cite{TheMEG:2016wtm}, $\mu\to 3e$~\cite{Bellgardt:1987du} and coherent $\mu\to e$ conversion in nuclei~\cite{Bertl:2006up}.

In addition, systematic efforts have also been devoted to understanding the theoretical aspects related to the interpretation of the absence of charged LFV (cLFV) signals in terms of limits on BSM physics. Background calculations were carried out for the so-called ``radiative''~\cite{Fael:2015gua,Fael:2016hnz,Pruna:2017upz} and ``rare''~\cite{Flores-Tlalpa:2015vga,Fael:2016yle,Pruna:2016spf} decays, \emph{i.e.} for the processes $l\to l'\gamma+2\nu$ and $l\to 3l'+2\nu$ (or $l\to l'+2l''+2\nu$), respectively.
Both these channels are important for the determination of the limits on the branching ratios of the two cLFV processes $l\to l'\gamma$ and $l\to 3l'$ (or $l\to l'+2l''$) because they provide an identical signal in the circumstance where invisible energy tends to zero, especially in view of the new experimental plans to improve the exploring power on these channels by MEG~II~\cite{Baldini:2013ke} and Mu3e~\cite{Blondel:2013ia}.
As for the phenomenological interpretation of the absence of a signal in terms of limits on the parameter space of potential BSM scenarios, a systematic EFT treatment was proposed~\cite{Pruna:2014asa,Pruna:2015jhf} and further developed with a particular focus on the muonic coherent LFV transitions~\cite{Crivellin:2016ebg,Crivellin:2017rmk}.
\vspace{-0.3cm}
\subsection*{The Muon $g-2$ experiment}
\noindent
A vast theoretical literature exists on the muonic anomalous magnetic moment $a_\mu$ and the current tension of $\sim 3$ standard deviations between the value obtained by the E821 experiment performed at the Brookhaven National Laboratory~\cite{Bennett:2006fi} and the theoretical prediction~\cite{Hagiwara:2011af}, and reviewing it in detail goes beyond the scope of this note. Instead, the attention of the reader is brought to the latest contributions and new phenomenological developments. Concerning the former, we recall that the relevant contributions can be classified in the following way~\cite{Hagiwara:2017zod}: quantum electrodynamics contributions~\cite{Aoyama:2012wk}, EW contributions~\cite{Gnendiger:2013pva}, hadronic light-by-light leading~\cite{Jegerlehner:2009ry} and next-to-leading~\cite{Colangelo:2014qya} order contributions, and hadronic vacuum polarisation leading~\cite{Davier:2010nc}, next-to-leading (see~\cite{Jegerlehner:2009ry} for a detailed review) and next-to-next-to-leading~\cite{Kurz:2014wya} order contributions.

Recently, a novel proposal~\cite{Calame:2015fva,Abbiendi:2016xup} suggested extracting the hadronic vacuum polarisation from the contribution to the muonic $g-2$ by exploiting $\mu e$ scattering.


{\par\medskip{\centering$\ast\ \ast\ \ast$\par}\medskip}
\let\stars\acknowledgments
\let\solong\acknowledgments
\noindent
This work was supported by the SNSF under contract 200021\_160156.


\end{document}